# PartDiffuser: Part-wise 3D Mesh Generation via Discrete Diffusion

### v1

@@ 2.2 Part-level 3D Generation
Another category of methods leverages Diffusion Transformer (DiT) models to achieve part-level generation, often bypassing the need for explicit segmentation results. Both PartCrafter [20] and PartPacker [33], for example, utilize multi-instance DiTs to automatically produce components. PartCrafter can operate from a single image to jointly form multiple, semantically distinct parts, enabling an end-to-end, part-conscious 3D mesh synthesis. PartPacker extends this concept by introducing a dual-volume DiT, which improves efficiency by modeling complementary spatial volumes. Frankenstein [39] employs a comparable strategy, using a VAE to consolidate multiple SDFs within a latent triplane space. Additionally, AutoPartGen [4] uses a latent diffusion model to sequentially generate parts in an autoregressive fashion, though this method is noted for being computationally intensive and offering limited user control. However, these methods often rely on continuous latent spaces or intermediate representations, leading to inevitable information loss during conversion to final meshes and difficulty in maintaining precise topological consistency across generated parts.
@@ 3.3.2 Part-Aware Diffusion Block
$Z_{\text{out}}=\text{CrossAttn}(Q=\text{LN}(Z^{\prime}),K=C_{\text{dyn}},V=C_{\text{dyn}})+Z^{\prime}$
where $C_{\text{dyn}}$ is the dynamically selected conditioning feature, constructed by concatenating $C_{\text{global}}$ with $C_{\text{part}_{i}}$ corresponding to the semantic part being processed.
Figure 3: Visualization of the composite attention mask during the parallel training phase, using $N=3$ parts as an example. This mask governs the attention mask described in Section 3.3.2, managing interactions across the $N$ noisy blocks, denoted $X_{t}$, and the $N$ clean blocks, denoted $X_{0}$. The legend details the four distinct attention behaviors: standard allowance for mesh tokens, specific allowance for padding tokens, and blockage enforced by either the Block Diffusion Mask or the Block-Aware Padding Mask.
Crucially, the self-attention module employs a specialized composite mask as an important tool to achieve part-wise generation. This mask combines two distinct components via logical intersection. The Block Diffusion Mask stemming from BD-LM [1] enables fine-grained control over the diffusion process by regulating the flow of information between different token blocks, each corresponding to a semantic part. It implements a hybrid attention structure: it permits full, bidirectional attention within each token block, while enforcing an autoregressive attention between token blocks, allowing part $i$ to attend only to previous parts $X_{<i}$. The second, Block-Aware Padding Mask, works in tight coordination with this hybrid structure. It supports the bidirectional intra-block attention by permitting full visibility within each token block, including padding tokens. Conversely, for the autoregressive inter-block attention, it strictly prevents attention to any padding tokens in other blocks, ensuring context is derived only from valid data. The example of the composite mask used during the training phase can be found in Figure 3.
@@ 3.3.4 Training and Sampling Strategy
Our training strategy is designed to efficiently teach the model the association between token blocks and their corresponding part-level geometric features. During the training phase, the model operates in a parallel mode, receiving the complete token sequence at once. Following [1], this sequence is logically divided into $2N$ blocks ($N$ blocks for the noisy version $X_{t}$ and $N$ blocks for the clean version $X_{0}$). The composite attention mask applied to this parallel block structure is visualized in Figure 3. For every token block in the sequence (whether from $X_{t}$ or $X_{0}$), the Key and Value for its corresponding cross-attention layer are statically set to its specific geometric condition $C_{\text{global}},C_{\text{part}_{i}}]$. This parallelization strategy allows the model to efficiently learn the precise correspondence between token sequences and specific part geometries.
In contrast, the generation process, as shown in Algorithm 1, is semi-autoregressive, generating one semantic part at a time. The sampling process is divided into $N$ strides, where $N$ is the total number of semantic parts determined by the input point cloud condition $C_{\text{pc}}$. When generating the $i$-th part, the sampler passes the index $i$ of the block currently being generated to the model. At this point, the internal Part-Aware Diffusion Block switches to sampling mode, and its cross-attention layer’s Key and Value are dynamically constructed as $C_{\text{global}},C_{\text{part}_{i}}]$. This dynamic conditioning mechanism forces the model, while denoising the current token block, to focus its attention exclusively on the global geometric feature and the specific geometric feature of the $i$-th part. Once the tokens for this block are fully denoised, the sampler proceeds to stride $i+1$, and the cross-attention condition switches to $C_{\text{global}},C_{\text{part}_{i+1}}]$. This process repeats until all $N$ parts have been generated.
Algorithm 1 Sampling Strategy
3:for $i\leftarrow 1$ to NN do
4:  $Z\leftarrow\text{SamplePrior}(B,L_{\text{block}})$
5:  $X_{\text{context}}\leftarrow\text{concat}(X_{\text{accum}},Z)$
7:  for $t\leftarrow T$ down to 11 do
8:    // Pass full context, focus on last block
10:    // CrossAttn uses $C_{\text{dyn}}$ due to block_idx
13:  end for
15:end for
16:return $X_{\text{accum}}$
@@ 4.1 Implementation Details
Our model is built upon a discrete diffusion framework similar to BD-LM[1], which we adapt to incorporate the proposed hierarchical point cloud features 3.3. The model comprises 0.3B parameters. We constrain the maximum sequence length to 4096 tokens, and the semi-autoregressive block size is set to 1024 tokens, aligning with our dataset’s filtering criteria. Training was conducted for two weeks utilizing 4 NVIDIA H100 80GB GPUs. We employed a cosine learning rate scheduler, annealing the learning rate from a peak of $3\times 10^{-4}$ down to $1\times 10^{-5}$ over the training duration.
@@ A Dataset Construction
B Sampling Analysis
B.1 Sampling Process
To intuitively illustrate the core mechanism of PartDiffuser, we visualize the step-by-step sampling process in Figure 7. Crucially, the visualization highlights a distinct behavior compared to conventional fully autoregressive (AR) approaches. In standard AR methods, mesh faces are typically generated sequentially along a fixed topological trajectory or backbone. In contrast, the intra-part diffusion process exhibits a parallel generation pattern: faces do not propagate linearly from a seed point. Instead, valid geometry materializes simultaneously across disparate spatial locations within the part’s volume. As observed in the intermediate steps, the model does not merely extend a continuous partial surface; rather, it often constructs a sparse structural ”framework” or ”skeleton” of the part first. This aligns with the confidence-based nature of discrete diffusion, where the model prioritizes high-confidence structural tokens. The model’s ability to establish the overall geometric scaffolding may help mitigate the risk of local over-complexity.
Besides, the visualization also verifies semi-autoregressive design, where subsequent parts are explicitly conditioned on the completed geometry of their predecessors. It can be observed that in the noisy stages, the newly emerging faces tend to align with the boundaries of the previously generated fixed parts. This confirms that the model utilizes the context from previous parts to ensure topological connectivity.
@@ B.1 Sampling Process
B.2 Efficiency Analysis
Impact of Parallel Generation.
In the experiments described above, we consistently utilized a high-confidence top-$k$ sampling strategy, strictly maintaining $k=1$, to ensure the optimal quality of the generated results. However, as mentioned in the conclusion of the main text and widely discussed in recent studies [42, 27], the capability for parallel generation represents a distinct advantage of discrete diffusion models over AR counterparts. Consequently, in this section, we conduct an additional brief experiment to investigate the trade-offs involved in this mechanism and explore how parallel generation configurations influence the quality of the generated meshes.
We analyze the trade-off between inference efficiency and generation quality by adjusting the parallel generation factor $k$, which sets the number of tokens unmasked at each diffusion step. Given that the block size $L$ for each part is normalized to 1024 tokens, the total number of sampling steps $T$ is calculated as $T=L/k$. We compare the standard sampling strategy with $k=1$ used in our main experiments against accelerated parallel settings with $k=2$ and $k=4$.
Table 3: Ablation study on Parallel Generation Factor $k$. We evaluate the impact of increasing the parallel token prediction count $k$ on the Objaverse dataset. $T$ denotes the total sampling steps per part. Time represents the average inference latency per mesh in seconds.
Top-$k$ Steps ($T$) Time (s) $\downarrow$ CD $\times 10^{3}$ $\downarrow$ HD $\downarrow$
1 (Default) 1024 63.8 17.813 0.238
2 512 33.7 27.657 0.299
4 256 17.1 44.720 0.321
Table 3 quantitatively illustrates the trade-off between inference efficiency and generation quality. While the default setting ($k=1$) serves as the high-fidelity baseline, increasing the parallel generation factor reveals significant potential for acceleration. Specifically, by elevating $k$ to $4$, we achieve a substantial $3.7\times$ speedup, drastically reducing the inference time to $17.1$s. However, this aggressive parallelization comes at the cost of geometric precision, as indicated by the increase in CD to $44.720$. As visualized in Figure 8, while the global topology remains consistent, increasing $k$ leads to increasingly irregular or erroneous high-frequency surface details. Therefore, these results demonstrate that to obtain higher-quality models, we perhaps should appropriately sacrifice efficiency requirements.
Figure 8: Visual comparison of varying $k$.

### v2

@@ 2.2 Part-level 3D Generation
Another category of methods leverages Diffusion Transformer (DiT) models to achieve part-level generation, often bypassing the need for explicit segmentation results. Both PartCrafter [20] and PartPacker [33], for example, utilize multi-instance DiTs to automatically produce components. PartCrafter can operate from a single image to jointly form multiple, semantically distinct parts, enabling an end-to-end, part-conscious 3D mesh synthesis. PartPacker extends this concept by introducing a dual-volume DiT, which improves efficiency by modeling complementary spatial volumes. Frankenstein [39] employs a comparable strategy, using a VAE to consolidate multiple SDFs within a latent triplane space. Additionally, AutoPartGen [4] uses a latent diffusion model to sequentially generate parts in an autoregressive fashion, though this method is noted for being computationally intensive and offering limited user control. Furthermore, BANG [44] dynamically divides complete 3D assets into individual parts through generative exploded dynamics. However, these methods often rely on continuous latent spaces or intermediate representations, leading to inevitable information loss during conversion to final meshes and difficulty in maintaining precise topological consistency across generated parts.
@@ 3.3.2 Part-Aware Diffusion Block
$\hat{Z}=\text{CrossAttn}(Q=\text{LN}(Z^{\prime}),K=C_{\text{dyn}},V=C_{\text{dyn}})+Z^{\prime}$
where $C_{\text{dyn}}$ is the dynamically selected conditioning feature, constructed by concatenating $C_{\text{global}}$ with $C_{\text{part}_{i}}$ corresponding to the semantic part being processed. Finally, $\hat{Z}$ is processed by a standard Feed-Forward Network with a residual connection to produce the ultimate block output $Z_{\text{out}}$.
The self-attention module employs a specialized composite mask as an important tool to achieve part-wise generation. Building upon block-wise semi-autoregressive principles [1], it permits full, bidirectional attention within each token block, while enforcing an autoregressive attention between token blocks, allowing part $i$ to attend only to previous parts $X_{<i}$. Furthermore, a Block-Aware Padding Mask works in coordination to manage padding token visibility across these modes. We defer the detailed visualization and formulation of this composite mask to the Appendix.
@@ 3.3.4 Training and Sampling Strategy
Following BD-LM [1], our model operates in a parallel mode during training, receiving the complete token sequence logically divided into noisy and clean blocks to efficiently teach the association between token blocks and part-level geometric features. For every token block, the cross-attention layers are conditioned on its specific geometric feature $C_{\text{global}},C_{\text{part}_{i}}]$.
During inference, the generation process is semi-autoregressive, reconstructing one semantic part at a time over $N$ strides. When generating the $i$-th part, we employ the sampling strategy from LLaDA [27] fro intra-block generation, iteratively predicting and selectively remasking tokens. The cross-attention condition is dynamically set to $C_{\text{global}},C_{\text{part}_{i}}]$, guiding the denoising process with both global and local geometries. Once fully denoised, the condition updates for the subsequent part, repeating until the entire mesh is generated.
@@ 4.1 Implementation Details
Our model is built upon a discrete diffusion framework similar to BD-LM[1], which we adapt to incorporate the proposed hierarchical point cloud features 3.3. The model comprises 0.3B parameters. We constrain the maximum sequence length to 4096 tokens, and the semi-autoregressive block size is set to 1024 tokens, aligning with our dataset’s filtering criteria. The training process was divided into two stages: pre-training on 8 NVIDIA H100 80GB GPUs for 3 days, followed by fine-tuning on 4 NVIDIA H100 80GB GPUs for two weeks.
@@ 4.4 Ablation Study
4.5 Efficiency Analysis
Following the sampling strategy of LLaDA [27], we investigate the potential for parallel generation—a key advantage of discrete diffusion over AR models. While our primary experiments utilize a high-fidelity strategy where the sampling budget $T$ per block is initially set to align with the block length, we further investigate the efficiency potential of our framework.
Table 3: Ablation study on Acceleration Factor $k$. We evaluate the impact of increasing $k$ on the Objaverse dataset. $T$ denotes the sampling steps for each block. Time represents the average inference latency per mesh in seconds.
Accel. ($k$) Steps ($T$) Time (s) $\downarrow$ CD $\times 10^{3}$ $\downarrow$ HD $\downarrow$
1 (Default) 1024 63.8 17.813 0.238
2 512 33.7 27.657 0.299
4 256 17.1 44.720 0.321
Table 3 quantitatively illustrates the trade-off between inference efficiency and generation quality. We introduce $k$ as the sampling acceleration factor, which controls the density of token transition in each step by proportionally reducing the diffusion steps $T$ for each block. While the default setting ($k=1,T=1024$) serves as the high-fidelity baseline, increasing $k$ to $4$ reduces the sampling budget to $T=256$. This achieves a substantial $3.7\times$ speedup, reducing the inference time to $17.1$s. However, this aggressive acceleration forces the model to transition more tokens within fewer iterations, often including tokens with lower prediction confidence. As indicated by the increase in CD to $44.720$, this trade-off comes at the cost of geometric precision. As visualized in Figure 5, while the global topology remains consistent, increasing $k$ leads to increasingly irregular or erroneous high-frequency surface details, such as irregular patches or fractured surfaces. These results demonstrate that to obtain high-fidelity results, an appropriate sampling budget must be maintained at the expense of inference speed.
Figure 5: Visual comparison of varying Acceleration Factor $k$.
@@ 5.1 Limitations and Future Work
6 Acknowledgments
The work was supported by the National Natural Science Foundation of China (No. 62406298).
@@ A Dataset Construction
B Training and Sampling Details
B.1 Training Strategy
Figure 7: Visualization of the composite attention mask during the parallel training phase, using $N=3$ parts as an example. This mask governs the attention mask used in Section 3.3.4, managing interactions across the $N$ noisy blocks, denoted $X_{t}$, and the $N$ clean blocks, denoted $X_{0}$. The legend details the four distinct attention behaviors: standard allowance for mesh tokens, specific allowance for padding tokens, and blockage enforced by either the Block Diffusion Mask or the Block-Aware Padding Mask.
As introduced in the main text, the self-attention module employs a specialized composite mask. This mask combines two distinct components via logical intersection. The Block Diffusion Mask stemming from BD-LM [1] enables fine-grained control over the diffusion process by regulating the flow of information between different blocks. The second component, the Block-Aware Padding Mask, works in tight coordination with this hybrid structure. It supports the bidirectional intra-block attention by permitting full visibility within the current denoising block, including padding tokens. Conversely, for the autoregressive inter-block attention, it strictly prevents attention to any padding tokens in other blocks. An example of the composite mask used during the training phase is illustrated in Figure 7.
During the training phase, the model operates in a parallel mode, receiving the complete token sequence at once. This sequence is logically divided into $2N$ blocks ($N$ blocks for the noisy version $X_{t}$ and $N$ blocks for the clean version $X_{0}$). The composite attention mask is applied to this parallel block structure. This parallelization strategy allows the model to efficiently learn the precise correspondence between token sequences and specific part geometries.
To balance training stability with generative coverage, we adopt a two-stage training curriculum. In the pre-training stage, we adopt the clipped noise schedule $\mathcal{U}[\beta,\omega]$ [1] to ensure efficient acquisition of core geometric features. In the fine-tuning stage, the model is fine-tuned under a full linear schedule $\mathcal{U}[0,1]$, enabling it to adapt to the complete denoising trajectory required during inference and improving overall robustness across all noise levels.
B.2 Sampling Strategy
Algorithm 1 Sampling Strategy
3:for $i\leftarrow 1$ to NN do
4:  $Z\leftarrow\text{SamplePrior}(B,L_{\text{block}})$
5:  $X_{\text{context}}\leftarrow\text{concat}(X_{\text{accum}},Z)$
7:  for $t\leftarrow T$ down to 11 do
8:    // Pass full context, focus on last block
10:    // CrossAttn uses $C_{\text{dyn}}$ due to block_idx
13:  end for
15:end for
16:return $X_{\text{accum}}$
Algorithm 1 provides the pseudo-code for our semi-autoregressive sampling process. The generation reconstructs one semantic part at a time, consisting of $N$ strides corresponding to the total number of parts determined by the input condition $C_{\text{pc}}$.
C Sampling Analysis
C.1 Robustness Analysis
To analyze robustness, we conducted a brief experiment using SAMPart3D [42] as an alternative segmentation upstream.
Figure 8: Experiment on robustness to different segmentation strategies.
Quantitatively, using SAMPart3D on Objaverse yields a CD of $19.11\times 10^{-3}$, closely matching the original $17.81$ using PartField. The analysis demonstrates that our model is reasonably robust to different segmentation strategies. The denoising process prioritizes the completeness and smoothness of local point cloud geometry rather than strict semantic boundaries. However, as shown in the comparison, different segmentation strategies can affect local detail representation. Coarser segments may reach this capacity sooner, leading to relative simplification, while more balanced partitions utilize the token budget effectively for higher fidelity.
C.2 Sampling Process
@@
To intuitively illustrate the core mechanism of PartDiffuser, we visualize the part-by-part sampling process in Figure 9. Crucially, the visualization highlights a distinct behavior compared to conventional fully AR approaches. In AR methods, mesh faces are typically generated sequentially along a fixed topological trajectory or backbone. In contrast, the intra-part diffusion process exhibits a parallel generation pattern: faces do not propagate linearly from a seed point. Instead, valid geometry materializes simultaneously across disparate spatial locations within the part’s volume. As observed in the intermediate steps, the model does not merely extend a continuous partial surface; rather, it often constructs a sparse structural ”framework” or ”skeleton” of the part first. This aligns with the confidence-based nature of discrete diffusion, where the model prioritizes high-confidence structural tokens. The model’s ability to establish the overall geometric scaffolding may help mitigate the risk of local over-complexity.
Besides, the visualization also verifies semi-autoregressive design, where subsequent parts are explicitly conditioned on the completed geometry of their predecessors. It can be observed that in the noisy stages, the newly emerging faces tend to align with the boundaries of the previously generated fixed parts. This may demonstrate that the model utilizes the context from previous parts to ensure topological connectivity.